OpenStudy (anonymous):

simplify 0.4x^3+0.5x^3-0.2x^3=

HanAkoSolo (jamierox4ev3r):

0.4^3+0.5^3-0.2^3=0.64+0.125-0.008

HanAkoSolo (jamierox4ev3r):

0.64+0.125-0.008=0.757